Use a direct proof to show that the sum of two odd integers is even

1
Expert's answer
2022-06-15T14:55:31-0400

If the integer "n" is odd, then "\exist k\in \Z" such that "n=2k+1."

The sum of two odd integers we can write as


"n_1+n_2=2k_1+1+2k_2+1"

"=2(k_1+k_2+1), k_1, k_2\in \Z"

Let "m=k_1+k_2+1, k_1, k_2\in \Z." Then "m\in \Z."

We see that "\exist m\in \Z" such that "n_1+n_2=2m."

This means that the sum "n_1+n_2" is an even number.

Therefore the sum of two odd integers is even.